OpenStudy (anonymous):

A credit card calculates interest using the average daily balance method. The card charges 23.4% annual interest rate on the average daily balance. The following transactions occurred during the March 1 – March 31 billing period. The average daily balance for the billing period is $6,149.68. Find the interest to be paid on April 1, the next billing date. Round to the nearest cent.

OpenStudy (anonymous):

is it 1439.03? 6149.68 x .234 i don't think that's right

OpenStudy (ranga):

Here is the compound interest formula: \[\Large A = P(1 + \frac{ r }{ n })^{nt}\] A = Account Balance P = Principal Amount r = Annual interest rate in decimal n = compounding period (compounded how many times a year) t = years invested ( here t will be in days because of average daily balance)

OpenStudy (ranga):

^^^ I think t should still be in years. The n should take care of daily compounding. P = 6,149.68 r = 0.234 n = 365 (due to daily compounding) t = 31 days (March 1 – March 31). t = 31/365 years Find A. Subtract P to get the interest.

OpenStudy (anonymous):

140.32?

OpenStudy (ranga):

I am getting $123.40 using Google calculator.

OpenStudy (anonymous):

so the A is the account balance for the month?

OpenStudy (ranga):

When interest is added to the credit card balance at the end of the month, the new balance is A.

OpenStudy (anonymous):

which would make it 6290

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

thank you

OpenStudy (ranga):

do you have the answer for the problem?